# S5: Scalable Semi-Supervised Semantic Segmentation in Remote Sensing

### v1

@@ Abstract
Semi-supervised semantic segmentation (S4) has advanced remote sensing (RS) analysis by leveraging unlabeled data through pseudo-labeling and consistency learning. However, existing S4 studies often rely on small-scale datasets and models, limiting their practical applicability. To address this, we propose S5, the first scalable framework for semi-supervised semantic segmentation in RS, which unlocks the potential of vast unlabeled Earth observation data typically underutilized due to costly pixel-level annotations. Built upon existing large-scale RS datasets, S5 introduces a data selection strategy that integrates entropy-based filtering and diversity expansion, resulting in the RS4P-1M dataset. Using this dataset, we systematically scales S4 methods by pre-training RS foundation models (RSFMs) of varying sizes on this extensive corpus, significantly boosting their performance on land cover segmentation and object detection tasks. Furthermore, during fine-tuning, we incorporate a Mixture-of-Experts (MoE)-based multi-dataset fine-tuning approach, which enables efficient adaptation to multiple RS benchmarks with fewer parameters. This approach improves the generalization and versatility of RSFMs across diverse RS benchmarks. The resulting RSFMs achieve state-of-the-art performance across all benchmarks, underscoring the viability of scaling semi-supervised learning for RS applications. All datasets, code, and models will be released at S5.
Figure 1: (a) Traditional S4 workflow: splitting the dataset into labeled and unlabeled subsets to improve model performance with few labeled samples. (b) The proposed S5 workflow: perform semi-supervised segmentation pretraining on both labeled and large-scale unlabeled datasets, followed by fine-tuning on RS benchmarks. (c) Comparison of segmentation performance across six RS segmentation benchmarks.
@@ 1 Introduction
Remote sensing (RS) semantic segmentation is a key task in RS image understanding, aiming to accurately classify each pixel to enable automatic recognition and analysis of land cover information [57]. However, traditional fully supervised segmentation methods heavily rely on pixel-level annotations, making the acquisition of high-quality training samples extremely costly [61]. This dependency also limits model generalization in diverse scenarios. To reduce the burden of manual labeling and lower costs, semi-supervised semantic segmentation (S4) [33] has gained increasing attention. S4 enhances RS image segmentation performance by combining a small amount of labeled images with a large number of unlabeled images during the training stage.
@@ 3.3 MoE-based Multiple Dataset Fine-tuning
Taking a standard Transformer block as an example, the computation of FFN-MoE begins with the the input feature $F^{\text{in}}$, it first passes through multi-head self-attention (MHSA), followed by a shared linear layer and a ReLU activation to obtain the intermediate feature:
@@ 4.2.1 Effectiveness of pre-training Dataset Curation
To improve the reliability of pseudo-labels and the diversity of the unlabeled data, we propose a data selection strategy based on entropy filtering and diversity expansion, resulting in a curated subset named MillionAID*. Experimental results show that MillionAID* achieves the best performance across all evaluation tasks: 67.66 on the iSAID validation set, outperforming both SAMRS (67.59) and MillionAID-random (66.32); and 79.77, 53.81, and 69.33 on Vaihingen, LoveDA, and DIOR-R, respectively. These results clearly demonstrate that the proposed method effectively selects high-quality and diverse unlabeled images. Compared with the existing SAMRS dataset, our method offers better scalability, thereby further enhancing the generalization and transferability of S4P models.
@@ 4.2.3 The Ratio of Shared and Specific Experts
Based on the experimental results in Table 3, we systematically evaluate the performance of different pre-training and fine-tuning strategies across four RS segmentation benchmarks, focusing on the advantages of MDF and the impact of the shared-to-specific expert ratio ($\alpha$) in the MoE-FFN design. Our analysis is based on the average accuracy (Average) across the four datasets.
First, under the SDF paradigm, S4P effectively improves the performance of the MAE backbone, achieving better results on all four downstream tasks, with the average accuracy increasing from 72.14 to 73.51. Furthermore, compared to SDF, MDF consistently outperforms across all four tasks, with the average accuracy rising from 73.51 to 73.75, while reducing the required number of parameters by nearly four times. This indicates that MDF helps the model learn more generalizable and robust feature representations under diverse data distributions. Next, we introduce MoE-FFN to enhance model capacity and examine the effect of different specific expert ratios $\alpha$ on performance. Specifically, as $\alpha$ increases from 1/8 to 1/4, performance steadily improves across multiple datasets, with average accuracy increasing from 73.80 to 74.15.
@@ 7.3 Implementation Details
The fine-tuning experiments are conducted on 4 NVIDIA RTX 4090 GPUs and 4 NVIDIA L20 GPUs. We evaluate the proposed S5 method using ViT-B, ViT-L, and ViT-H backbones on four RS semantic segmentation benchmarks and two RS object detection datasets. Detailed training configurations are provided in Table 4.

### v2

@@ Abstract
Semi-supervised semantic segmentation (S4) has advanced remote sensing (RS) analysis by leveraging unlabeled data through pseudo-labeling and consistency learning. However, existing S4 studies often rely on small-scale datasets and models, limiting their practical applicability. To address this, we propose S5, the first scalable framework for semi-supervised semantic segmentation in RS, which unlocks the potential of vast unlabeled Earth observation data typically underutilized due to costly pixel-level annotations. Built upon existing large-scale RS datasets, S5 introduces a data selection strategy that integrates entropy-based filtering and diversity expansion, resulting in the RS4P-1M dataset. Using this dataset, we systematically scale up S4 methods by pre-training RS foundation models (RSFMs) of varying sizes on this extensive corpus, significantly boosting their performance on land cover segmentation and object detection tasks. Furthermore, during fine-tuning, we incorporate a Mixture-of-Experts (MoE)-based multi-dataset fine-tuning approach, which enables efficient adaptation to multiple RS benchmarks with fewer parameters. This approach improves the generalization and versatility of RSFMs across diverse RS benchmarks. The resulting RSFMs achieve state-of-the-art performance across all benchmarks, underscoring the viability of scaling semi-supervised learning for RS applications. All datasets, code, and models will be released at S5.
Figure 1: (a) Traditional S4 workflow: splitting the dataset into labeled and unlabeled subsets to improve model performance with few labeled samples. (b) The proposed S5 workflow: perform semi-supervised segmentation pretraining on both labeled and large-scale unlabeled datasets, followed by fine-tuning on RS benchmarks. (c) Comparison of performance across four RS segmentation and two object detection benchmarks.
@@ 1 Introduction
Remote sensing (RS) semantic segmentation is a key task in RS image understanding, aiming to accurately classify each pixel to enable automatic recognition and analysis of land cover information [57]. However, traditional fully supervised segmentation methods heavily rely on pixel-level annotations, making the acquisition of high-quality training samples extremely costly. This dependency also limits model generalization in diverse scenarios. To reduce the burden of manual labeling and lower costs, semi-supervised semantic segmentation (S4) [33] has gained increasing attention. S4 enhances RS image segmentation performance by combining a small amount of labeled images with a large number of unlabeled images during the training stage.
@@ 3.3 MoE-based Multiple Dataset Fine-tuning
Taking a standard Transformer block as an example, the computation of FFN-MoE begins with the input feature $F^{\text{in}}$, it first passes through multi-head self-attention (MHSA), followed by a shared linear layer and a ReLU activation to obtain the intermediate feature:
@@ 4.2.1 Effectiveness of pre-training Dataset Curation
To improve the reliability of pseudo-labels and the diversity of the unlabeled data, we propose a data selection strategy based on entropy filtering and diversity expansion, resulting in a curated subset named MillionAID*. Experimental results show that MillionAID* achieves the best performance across all evaluation tasks: 67.66 on the iSAID validation set, outperforming both SAMRS (67.59) and MillionAID-random (66.32); and 79.77, 53.81, and 69.65 on Vaihingen, LoveDA, and DIOR-R, respectively. These results clearly demonstrate that the proposed method effectively selects high-quality and diverse unlabeled images. Compared with the existing SAMRS dataset, our method offers better scalability, thereby further enhancing the generalization and transferability of S4P models.
@@ 4.2.3 The Ratio of Shared and Specific Experts
Based on the results in Table 3, we evaluate various pre-training and fine-tuning strategies across four RS segmentation benchmarks, highlighting the advantages of MDF and the effect of the shared-to-specific expert ratio ($\alpha$) in the MoE-FFN design. The analysis is conducted using the average accuracy (Average) across the four datasets.
First, under the SDF paradigm, S4P effectively enhances the MAE backbone, improving performance across all four downstream tasks with the average accuracy rising from 72.14 to 73.51. Compared with SDF, MDF achieves consistently better results on all tasks, increasing the average accuracy to 73.75 while reducing parameters by nearly four times. This demonstrates that MDF enables the model to learn more general and robust feature representations under diverse data distributions. Finally, we introduce MoE-FFN to expand model capacity and analyze the impact of the specific expert ratio ($\alpha$). As $\alpha$ increases from 1/8 to 1/4, the average accuracy steadily improves from 73.80 to 74.15 across multiple datasets.
@@ 7.3 Implementation Details
The fine-tuning experiments are conducted on 4 NVIDIA RTX 4090 GPUs and 4 NVIDIA L20 GPUs. We evaluate the proposed S5 method using ViT-B, ViT-L, and ViT-H backbones on four RS semantic segmentation benchmarks and two RS object detection datasets. Detailed training configurations are provided in Table 4. For ease of comparison, all parameter counts are measured at an input resolution of $512\times 512$.